OpenStudy (anonymous):

What is the solution set (ordered pairs) for this inequality? y ≤ x + 2

OpenStudy (amistre64):

first graph the boundary line y = x+2

OpenStudy (amistre64):

then pick a point that is clearly NOT on the line; like 0,0 and see if its good or bad if good, use that side, if bad, use the other side

OpenStudy (amistre64):



OpenStudy (amistre64):

since: 0 <= 0+2 is true; we shade the side the has (0,0) in it

OpenStudy (amistre64):



OpenStudy (amistre64):

i spose the x>= and y<= would satisfy the solution sets

OpenStudy (amistre64):

{(x,y): x>= -2, y <= 2} something like that

OpenStudy (anonymous):

I can choose from..(there ca be more than one answer..) (-1, 1) (0, 3) (2, 4) (3, 2)

OpenStudy (amistre64):

well, given options to choose from, its tends to be simpler to just plug them in and see what fits :)

OpenStudy (amistre64):

y <= x + 2 1 -1 ------------ 1 <= -1+2 is true y <= x + 2 3 0 ---------- 3 <= 0+2 is false y <= x + 2 4 2 ----------- 4 <= 2+2 is true y <= x + 2 2 3 ---------- 2 <= 3+2 is false

OpenStudy (anonymous):

You lost me there lol

OpenStudy (amistre64):

i just pluged in the values for (x,y) that they gave us to see what rang true

OpenStudy (anonymous):

Which ones were true?

OpenStudy (amistre64):

well, seeing that i tried them in order .... id say the ones that are marked true :)

OpenStudy (anonymous):

so (3,2) and (2,4) are true?

OpenStudy (amistre64):

(-1, 1) T (0, 3) F (2, 4) T (3, 2) F im pretty sure i did the math right ... you might wanna dbl chk it tho

OpenStudy (anonymous):

It was right thanks :)

OpenStudy (amistre64):

yay!! :) good luck